The base of a triangle is 11 cm more than its height. The area is 90cm^2. What is the measure of the height?
Answer by checkley77(12844)

AREA=BH/2
90=(H+11)H/2
90=(H^2+11H)/2
H^2+11H=90*2
H^2+11H-180=0
(H+20)(H-9)=0
H-9=0
H=9 ANS. FOR THE HEIGHT.
9+11=20 ANS. FOR THE BASE.
PROOF:
90=9*20/2
90=180/2
90=90